### v2

# Anisotropic collective excitations of Bose gases in modified Newtonian dynamics

Abstract
Collective excitations are fundamental in quantum many-body physics, yet their spectra have traditionally been studied within Newtonian dynamics. In this paper, we investigate collective excitations in Bose gases under Modified Newtonian Dynamics (MOND). We derive an anisotropic excitation spectrum in the MOND regime. This anisotropy arises directly from the intrinsic nonlinear structure of the MOND Poisson equation, forming a distinctive signature of the modified gravitational response. We then analyze the Jeans instability, obtaining analytic expressions for the direction-dependent critical wavelength and mass. These results advance our understanding of collective behavior in quantum systems under modified dynamics and establish clear theoretical signatures for testing MOND-like effects in quantum simulators.
Introduction. Collective excitations represent a fundamental concept in quantum many-body physics. They characterize the intrinsic quantum nature of many-particle systems and provide key insights into emergent phenomena such as superfluidity and superconductivity Griffin (1993); Bruus and Flensberg (2004); Bloch et al. (2008); Coleman (2015). These low-energy modes describe how a quantum system responds to small perturbations. Their properties have traditionally been studied within the framework of Newtonian dynamics. In this work, we extend this paradigm by investigating collective excitations in a Bose gas under Modified Newtonian Dynamics (MOND). This exploration is motivated by the possibility of uncovering novel physical behavior in quantum many-body systems that fundamentally distinguishes MOND from its Newtonian counterpart.
MOND, proposed by Milgrom in 1983, posits a modification to the classical inertial or gravitational laws at accelerations below a critical scale $a_{0}\sim 10^{-10}\,\mathrm{m\,s}^{-2}$ Milgrom (1983). It provides a successful empirical description of galaxy rotation curves without invoking dark matter Bekenstein and Milgrom (1984); Garay et al. (2000); Bekenstein (2004); Famaey and McGaugh (2012); Milgrom (2014); Bugg (2015); Koyama (2016). The theory is inherently nonlinear and, in its deep-MOND regime, exhibits scale invariance Milgrom (2025), leading to distinct gravitational responses and scaling laws absent in Newtonian gravity.
Direct tests of MOND in terrestrial laboratories are challenging due to the extremely low acceleration scale $a_{0}$. However, quantum simulators, particularly ultracold atomic Bose-Einstein condensates (BECs), offer a versatile platform to emulate features of gravitational physics in a controlled setting Dalfovo et al. (1999); Fedichev and Fischer (2003, 2004); Boehmer and Harko (2007); Pethick and Smith (2008); Lewenstein et al. (2012); Steinhauer (2014); Das and Bhaduri (2015); Steinhauer (2016); Gross and Bloch (2017); Chä and Fischer (2017); Eckel et al. (2018); Barceló (2019); Muñoz de Nova et al. (2019); Kolobov et al. (2021); Holanda Ribeiro and Fischer (2023); Pal and Fischer (2024); Haddad (2025). A concrete proposal for mimicking a Newtonian potential was put forward by O’Dell et al. O’dell et al. (2000), who showed that six appropriately arranged off-resonant laser beams can induce an attractive $1/r$ interaction between neutral atoms. In the near-zone limit, the faster-decaying $1/r^{3}$ dipole-dipole terms average out, leaving the desired long-range $1/r$ potential. Although a self-bound condensate with such an interaction has not yet been realized experimentally, the required physical parameters have been analyzed in detail Giovanazzi et al. (2001), and the theoretical properties of the resulting system have been studied extensively Cartarius et al. (2008). Self-gravitating BECs themselves have been widely investigated as analog models for dark-matter halos Chavanis (2011, 2017, 2020, 2025). Their collective excitations in the Newtonian limit are described by the dispersion relation
$\omega^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+\frac{gn_{0}}{m}k^{2}-4\pi G\rho_{0},$ (1)
where the terms correspond to quantum pressure, short-range contact interactions, and gravitational attraction, respectively. Here, $n_{0}$ is the condensate density, $\rho_{0}=mn_{0}$ the mass density, $g$ the interaction strength, and $G$ Newton’s constant. Recent theoretical work has begun exploring ground-state properties of BECs under MOND-like logarithmic potentials, revealing distinctive scaling laws and significant cloud enlargement compared to Newtonian traps Liu (2026). These findings motivate a deeper investigation into dynamical properties, particularly the full collective excitation spectrum, within a consistent MOND framework.
We analyze collective modes of a self-gravitating Bose-Einstein condensate within the MOND framework. Beginning with the coupled Gross–Pitaevskii and MOND Poisson equations, a linear stability analysis via the Bogoliubov–de Gennes approach yields an anisotropic dispersion relation,
$\omega^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+\frac{gn_{0}}{m}k^{2}-\frac{4\pi G\rho_{0}k^{2}}{D(\mathbf{k},\theta)},$ (2)
where the factor $D(\mathbf{k},\theta)$ embodies the directional dependence emerging from the nonlinear MOND field equation. The anisotropy thus originates not from anisotropic interparticle interactions (such as dipolar coupling) but is intrinsic to the modified gravitational response. Consequently, the excitation frequency depends on both the wavenumber $k$ and the angle $\theta$ between the perturbation wavevector and the background gravitational field. This angular dependence makes the Jeans instability criterion directional: critical wavelengths and masses for gravitational collapse vary systematically with orientation. Our results establish a clear anisotropic signature of MOND, offering a novel pathway to test phenomena predicted by modified gravity/dynamics in quantum simulators with engineered long-range interactions.
Model. The system consists of a self-gravitating BEC described by a macroscopic wave function $\psi(\mathbf{r},t)$ and a gravitational potential $\Phi(\mathbf{r},t)$. Its dynamics follows from the Lagrangian density
$\mathcal{L}[\psi,\psi^{*},\Phi]=\mathcal{L}_{\mathrm{GP}}[\psi,\psi^{*}]+\mathcal{L}_{\mathrm{MOND}}[\psi,\psi^{*},\Phi],$ (3)
with the Gross–Pitaevskii (GP) part Pitaevskii and Stringari (2016)
$\mathcal{L}_{\mathrm{GP}}$ $=\frac{i\hbar}{2}\left(\psi^{*}\partial_{t}\psi-\psi\partial_{t}\psi^{*}\right)-\frac{\hbar^{2}}{2m}|\nabla\psi|^{2}$ (4)
$\quad-V_{\mathrm{ext}}|\psi|^{2}-\frac{g}{2}|\psi|^{4},$
and the MOND gravitational part Bekenstein and Milgrom (1984)
$\mathcal{L}_{\mathrm{MOND}}=-m\Phi|\psi|^{2}-\frac{a_{0}^{2}}{8\pi G}\,\mathcal{F}\!\left(\frac{|\nabla\Phi|^{2}}{a_{0}^{2}}\right).$ (5)
Here $m$ is the atomic mass, $V_{\mathrm{ext}}$ an external trapping potential, and $g=4\pi\hbar^{2}a_{s}/m$ the contact interaction strength, where $a_{s}$ is the $s$-wave scattering length. The MOND acceleration scale is denoted by $a_{0}$. The function $\mathcal{F}(x^{2})$ is related to the MOND interpolation function $\mu(x)$ via $\mu(x)=\mathcal{F}^{\prime}(x^{2})$. In the Newtonian limit $|\nabla\Phi|\gg a_{0}$, one has $\mu\to 1$, while in the deep-MOND limit $|\nabla\Phi|\ll a_{0}$, $\mu(x)\approx x$.
Varying the action $S=\int dt\int d^{3}\mathbf{r}\,\mathcal{L}$ with respect to $\psi^{*}$ and $\Phi$ yields the equations of motion. Variation with respect to $\psi^{*}$ gives the GP equation
$i\hbar\frac{\partial\psi}{\partial t}=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\mathrm{ext}}+g|\psi|^{2}+m\Phi\right]\psi,$ (6)
and variation with respect to $\Phi$ leads to the MOND Poisson equation
$\nabla\cdot\left[\,\mu\!\left(\frac{|\nabla\Phi|}{a_{0}}\right)\nabla\Phi\,\right]=4\pi G\,m\,|\psi|^{2}.$ (7)
Setting $\mu=1$ recovers the standard Poisson equation $\nabla^{2}\Phi=4\pi Gm|\psi|^{2}$; Eqs. (6) and (7) then reduce to the Gross–Pitaevskii–Poisson (or GP-Newton) system describing a self-gravitating BEC in Newtonian gravity.
Bogoliubov-de Gennes Theory. To study small-amplitude collective modes, we expand the wave function and gravitational potential around their equilibrium values:
$\psi(\mathbf{r},t)$ $=e^{-i\mu_{c}t/\hbar}\left[\psi_{0}+\delta\varphi(\mathbf{r},t)\right],$ (8)
$\Phi(\mathbf{r},t)$ $=\Phi_{0}(\mathbf{r})+\delta\Phi(\mathbf{r},t),$ (9)
with $|\delta\varphi|\ll|\psi_{0}|$ and $|\delta\Phi|\ll|\Phi_{0}|$. Here $\psi_{0}$ is the equilibrium condensate wave function, $\Phi_{0}$ the equilibrium gravitational potential, and $\mu_{c}$ the chemical potential. Substituting these expansions into Eqs. (6) and (7) and linearizing in the perturbations yields
$i\hbar\frac{\partial\delta\varphi}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\delta\varphi+gn_{0}(\delta\varphi+\delta\varphi^{*})+m\psi_{0}\delta\Phi,$ (10)
and
$\nabla\cdot\left[\mu_{0}\nabla\delta\Phi+\frac{\mu_{0}^{\prime}}{a_{0}}(\hat{e}\cdot\nabla\delta\Phi)\mathbf{g}_{0}\right]=4\pi Gm\psi_{0}(\delta\varphi+\delta\varphi^{*}).$ (11)
The equilibrium density is $n_{0}=|\psi_{0}|^{2}$. The background gravitational field $\mathbf{g}_{0}=-\nabla\Phi_{0}$ has magnitude $\mathcal{G}_{0}=|\mathbf{g}_{0}|$ and direction $\hat{e}=\mathbf{g}_{0}/\mathcal{G}_{0}$. The coefficients $\mu_{0}=\mu(\mathcal{G}_{0}/a_{0})$ and $\mu_{0}^{\prime}=d\mu/dx|_{x=\mathcal{G}_{0}/a_{0}}$ are the MOND interpolation function and its derivative evaluated at the scaled field strength.
We consider a locally homogeneous background and look for plane-wave perturbations
$\delta\varphi(\mathbf{r},t)$ $=ue^{i(\mathbf{k}\cdot\mathbf{r}-\omega t)}+v^{*}e^{-i(\mathbf{k}\cdot\mathbf{r}-\omega t)},$ (12)
$\delta\Phi(\mathbf{r},t)$ $=we^{i(\mathbf{k}\cdot\mathbf{r}-\omega t)}+w^{*}e^{-i(\mathbf{k}\cdot\mathbf{r}-\omega t)},$ (13)
with constant amplitudes $u$, $v$, $w$, wavevector $\mathbf{k}$, and frequency $\omega$. Substituting these forms into the linearized equations and matching Fourier coefficients leads to the Bogoliubov–de Gennes (BdG) equations Pethick and Smith (2008)
$\hbar\omega u$ $=\frac{\hbar^{2}k^{2}}{2m}u+gn_{0}(u+v)+m\psi_{0}w,$ (14)
$-\hbar\omega v$ $=\frac{\hbar^{2}k^{2}}{2m}v+gn_{0}(u+v)+m\psi_{0}w,$ (15)
together with the Fourier-space MOND equation
$\left[-\mu_{0}k^{2}-\frac{\mu_{0}^{\prime}\mathcal{G}_{0}}{a_{0}}(\mathbf{k}\cdot\hat{e})^{2}\right]w=4\pi Gm\psi_{0}(u+v).$ (16)
Defining the anisotropic denominator
$D(\mathbf{k},\theta)=\mu_{0}k^{2}+\frac{\mu_{0}^{\prime}\mathcal{G}_{0}}{a_{0}}(\mathbf{k}\cdot\hat{e})^{2},$ (17)
we solve Eq. (16) for $w$:
$w=-\frac{4\pi Gm\psi_{0}}{D(\mathbf{k},\theta)}(u+v).$ (18)
Eliminating $w$ from the BdG equations gives a $2\times 2$ eigenvalue problem. Its solvability condition yields the dispersion relation
$\omega^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+\frac{k^{2}}{m}\left[gn_{0}-\frac{4\pi Gm^{2}n_{0}}{D(\mathbf{k},\theta)}\right],$ (19)
which is the central result of our analysis. The factor $D(\mathbf{k},\theta)$ depends on $(\mathbf{k}\cdot\hat{e})^{2}$, introducing an anisotropy: the excitation frequency varies with the angle $\theta$ between $\mathbf{k}$ and the background field direction $\hat{e}$.
In the Newtonian limit ($\mathcal{G}_{0}\gg a_{0}$), $\mu_{0}\to 1$ and $\mu_{0}^{\prime}\to 0$, so $D(\mathbf{k},\theta)\to k^{2}$ and Eq. (19) reduces to the isotropic quantum Jeans dispersion
$\omega^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+\frac{gn_{0}}{m}k^{2}-4\pi Gmn_{0}.$ (20)
In the deep-MOND regime ($\mathcal{G}_{0}\ll a_{0}$), the interpolation function behaves as $\mu(x)\approx x$, giving $\mu_{0}\approx\mathcal{G}_{0}/a_{0}$ and $\mu_{0}^{\prime}\approx 1$. Then
$D(\mathbf{k},\theta)\approx\frac{\mathcal{G}_{0}}{a_{0}}\left[k^{2}+(\mathbf{k}\cdot\hat{e})^{2}\right]=\frac{\mathcal{G}_{0}}{a_{0}}k^{2}(1+\cos^{2}\theta),$ (21)
with $\cos\theta=\hat{k}\cdot\hat{e}$. Substituting into Eq. (19) gives the deep-MOND dispersion relation
$\omega^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+\frac{gn_{0}}{m}k^{2}-\frac{4\pi Gmn_{0}a_{0}}{\mathcal{G}_{0}(1+\cos^{2}\theta)}.$ (22)
The anisotropy is now explicit: the gravitational term depends on $\theta$. For $\theta=0$ (wavevector parallel to $\hat{e}$), the denominator is $2\mathcal{G}_{0}$, yielding the weakest gravitational correction; for $\theta=\pi/2$ (perpendicular case), the denominator is $\mathcal{G}_{0}$, giving a correction twice as strong.
In the limit of negligible gravity ($G\to 0$), Eq. (22) reduces to the familiar Bogoliubov excitation spectrum $\omega^{2}=\hbar^{2}k^{4}/(4m^{2})+(gn_{0}/m)k^{2}$. In the classical limit ($\hbar\to 0$), the quantum pressure term vanishes, giving the purely classical dispersion relation
$\omega^{2}=\frac{gn_{0}}{m}k^{2}-\frac{4\pi Gmn_{0}a_{0}}{\mathcal{G}_{0}(1+\cos^{2}\theta)}.$ (23)
This classical expression highlights the anisotropic gravitational contribution without quantum effects.
To highlight the scaling, we introduce the healing length $\xi=\hbar/\sqrt{2mgn_{0}}$ and the dimensionless quantities
$\tilde{k}=k\xi,\qquad\tilde{\omega}=\frac{\hbar\omega}{gn_{0}},\qquad\chi=\frac{4\pi Gm^{2}\xi^{2}}{g},\qquad\eta=\frac{\mathcal{G}_{0}}{a_{0}}.$ (24)
Here $\chi$ measures the relative strength of gravity, and $\eta\ll 1$ in the deep-MOND regime. Eq. (22) then becomes
$\tilde{\omega}^{2}=\tilde{k}^{4}+2\tilde{k}^{2}-\frac{2\chi}{\eta(1+\cos^{2}\theta)}.$ (25)
To illustrate the anisotropy, we first plot in Fig. 1 the dimensionless squared frequency $\tilde{\omega}^{2}$ as a function of the angle $\theta$ for several fixed dimensionless wavenumbers $\tilde{k}$. For each $\tilde{k}$, $\tilde{\omega}^{2}$ decreases monotonically from $\theta=0$ to $\theta=\pi/2$, confirming that perturbations perpendicular to the background gravitational field are more unstable—i.e., have a lower or more negative squared frequency—than parallel ones at the same wavenumber. Consequently, the critical wavenumber for instability (where $\tilde{\omega}^{2}=0$) is smaller for perpendicular perturbations and larger for parallel ones. This means that, to become unstable, modes parallel to the background field require a shorter wavelength (larger $k$) compared to perpendicular modes.
Figure 1: Squared frequency $\tilde{\omega}^{2}$ versus angle $\theta$ for fixed wavenumbers $\tilde{k}=0.5$ (dashed), $\tilde{k}=1.0$ (dot-dashed), and $\tilde{k}=1.5$ (solid), with $\chi=1$ and $\eta=0.1$.
The full anisotropic landscape is displayed in Fig. 2 as a contour plot of $\tilde{\omega}^{2}$ in the $\tilde{k}-$$\theta$ plane. The red dashed contour marks the stability boundary $\tilde{\omega}^{2}=0$, which separates the parameter space into two regions: the stable collective-excitation region on the right ($\tilde{\omega}^{2}>0$) and the unstable, collapse-prone region on the left ($\tilde{\omega}^{2}<0$). The shape of this boundary clearly shows that the critical wavenumber for stability decreases as $\theta$ increases from $\theta=0$ to $\theta=\pi/2$, confirming the directional dependence of the Jeans instability.
Figure 2: Contour plot of the squared frequency $\tilde{\omega}^{2}$ in the $\tilde{k}-$$\theta$ plane for $\chi=1$, $\eta=0.1$. The red dashed line indicates the stability boundary $\tilde{\omega}^{2}=0$, separating the stable (right) and unstable (left) regions.
Jeans Instability Analysis. In the long-wavelength limit $\tilde{k}\to 0$, Eq. (25) gives
$\lim_{\tilde{k}\to 0}\tilde{\omega}^{2}=-\frac{2\chi}{\eta(1+\cos^{2}\theta)}<0,$ (26)
indicating Jeans instability for all directions, with a growth rate that depends on $\theta$. The critical wavenumber $k_{J}$ at which $\omega=0$ satisfies
$\tilde{k}_{J}^{4}+2\tilde{k}_{J}^{2}=\frac{2\chi}{\eta(1+\cos^{2}\theta)},$ (27)
where $\tilde{k}_{J}=k_{J}\xi$ is the dimensionless critical wavenumber. Solving this quadratic equation for $\tilde{k}_{J}^{2}$ yields
$\tilde{k}_{J}^{2}=\sqrt{1+\frac{2\chi}{\eta(1+\cos^{2}\theta)}}-1,$ (28)
or equivalently
$k_{J}(\theta)=\frac{1}{\xi}\sqrt{\sqrt{1+\frac{2\chi}{\eta(1+\cos^{2}\theta)}}-1}.$ (29)
This is a decreasing function of $\cos^{2}\theta$; hence $k_{J}$ is largest for $\theta=\pi/2$ and smallest for $\theta=0$.
The Jeans wavelength, defined as $\lambda_{J}=2\pi/k_{J}$, becomes
$\lambda_{J}(\theta)=2\pi\xi\left[\sqrt{1+\frac{2\chi}{\eta(1+\cos^{2}\theta)}}-1\right]^{-1/2},$ (30)
which is smallest for $\theta=\pi/2$ and largest for $\theta=0$. This anisotropic Jeans criterion contrasts sharply with the isotropic one in Newtonian gravity.
The Jeans mass $M_{J}$, defined as the mass contained within a sphere of diameter $\lambda_{J}$, is given by
$M_{J}(\theta)=\frac{4\pi}{3}\rho_{0}\left(\frac{\lambda_{J}(\theta)}{2}\right)^{3}=\frac{\pi}{6}\rho_{0}\lambda_{J}^{3}(\theta),$ (31)
where $\rho_{0}=mn_{0}$ is the mass density. Substituting Eq. (30) gives
$M_{J}(\theta)=\frac{\pi}{6}\rho_{0}(2\pi\xi)^{3}\left[\sqrt{1+\frac{2\chi}{\eta(1+\cos^{2}\theta)}}-1\right]^{-3/2}.$ (32)
In the classical limit ($\hbar\to 0$), the quantum pressure term vanishes and the dispersion relation reduces to Eq. (23). Setting $\omega=0$ in that equation gives the classical Jeans wavenumber
$k_{J,\mathrm{class}}(\theta)=\sqrt{\frac{4\pi Gm^{2}a_{0}}{g\,\mathcal{G}_{0}\,(1+\cos^{2}\theta)}},$ (33)
which likewise exhibits a clear $\theta$-dependence. Hence, the anisotropy of the Jeans scale is generic consequence of the MOND nonlinearity, persisting in both quantum and classical regimes.
To illustrate the angular anisotropy of the collapse scale, we introduce the normalized Jeans mass $\widetilde{M}_{J}(\theta)=M_{J}(\theta)/M_{J}(0)$, which measures the mass ratio relative to the parallel direction ($\theta=0$). From Eq. (32), we have
$\widetilde{M}_{J}(\theta)=\left[\frac{\sqrt{1+\chi/\eta}-1}{\sqrt{1+\frac{2\chi}{\eta(1+\cos^{2}\theta)}}-1}\right]^{3/2}.$ (34)
This function decreases monotonically from $1$ at $\theta=0$ to its minimum at $\theta=\pi/2$, with $\widetilde{M}_{J}(\pi/2)=[(\sqrt{1+\chi/\eta}-1)/(\sqrt{1+2\chi/\eta}-1)]^{3/2}<1$. Fig. 3 displays $\widetilde{M}_{J}(\theta)$ in polar coordinates. The radial coordinate represents $\widetilde{M}_{J}(\theta)$; the dashed circle marks the isotropic Newtonian limit (normalized to $1$). The plot reveals a pronounced anisotropy: the Jeans mass is largest along the direction parallel to the background field ($\theta=0$ and $\pi$) and smallest in the perpendicular directions ($\theta=\pi/2$ and $3\pi/2$). The anisotropy factor $M_{J}(0)/M_{J}(\pi/2)\approx 2.5$ for these parameters, underscoring a strong directional dependence of the collapse scale in the MOND regime.
Figure 3: Polar plot of the normalized Jeans mass $\widetilde{M}_{J}(\theta)$. The radial coordinate is $\widetilde{M}_{J}(\theta)$; the dashed circle indicates the isotropic Newtonian limit (normalized to $1$). Parameters are the same as in Fig. 1.
The angular dependence of the Jeans mass has important implications for structure formation in self-gravitating systems under MOND. Unlike Newtonian gravity, where collapse occurs isotropically, MOND predicts that gravitational instability is more efficient in directions perpendicular to the background field, favoring the formation of anisotropic structures. On the other hand, the anisotropic instability can be a hallmark of MOND and provides a distinctive signature that could be tested in quantum simulators with engineered long-range interactions.
Summary and outlook. We have derived the collective excitation spectrum of a self-gravitating Bose-Einstein condensate in the MOND framework. The dispersion relation exhibits a clear anisotropy, depending on the angle between the perturbation wavevector and the background gravitational field. This anisotropy stems from the nonlinearity of the MOND Poisson equation and disappears in the Newtonian limit. In the deep-MOND regime, the Jeans instability becomes direction-dependent, with perpendicular perturbations being more unstable than parallel ones.
Our work extends the study of quantum fluids to the low-acceleration regime described by MOND. The anisotropic excitation spectrum offers a distinct signature that could distinguish MOND-like effects from standard Newtonian behavior. Although derived here for a BEC, the same anisotropy should appear in other quantum liquids and in classical systems, because its origin lies in the modified gravitational response. This universality makes the effect a promising target for experimental simulation.
Acknowledgments
This work was supported by the Open Fund of Key Laboratory of Multiscale Spin Physics (Ministry of Education), Beijing Normal University (Grant No. SPIN2024N03), and the Scientific Research Startup Foundation for High-Level Talents at Anqing Normal University (Grant No. 241042).